The product of two numbers is -27. One of the number is 12 less than the other. Which of the following could be the two numbers:
a) 9 and 3
b) 1 and -27
c) 9 and -3
d) -13 and -14
I was calculating b) 1 and -27.
Want to make sure my calculations are correct

Answer by ewatrrr(24785)

n(n-12) = -27
n^2 - 12n +27 = 0
(n-9)(n-3) = 0
n = 3, 9
(n-12) is -9, -3
9 and -3 one such pair